nickadai:

This was an exceptionally dry year for portions of the southwestern United States. Monthly precipitation in Phoenix, Arizona, was recorded in the table and is modeled by y = –0.04088x2 + 0.4485x + 1.862. In what month did Phoenix receive the lowest amount of precipitation? Month (x) Precipitation January 2.27 inches February ? March ? April ? May ? June ? July ? August ? September 2.59 inches October ? November ? December ? Sketch a graph or fill in the table to answer the question. January February November December

MaxTon:

home work or test q uestion

nickadai:

regular assigment

nickadai:

Hello?

Extrinix:

Can you screenshot the table please.

MaxTon:

okay well can you like show us the table

nickadai:

These are the answers: January February November December

nickadai:

its not january btw

nickadai:

?

MaxTon:

dang this might take me a little bit

nickadai:

Im just happy for any help

MaxTon:

well josh is typing witch means he has the answer so if i don't back in time he has it

nickadai:

hes been "typing" for over 6 min.

nickadai:

?

Extrinix:

So to start off, you should know that the precipitation is y and the month is x (as shown). Now, to solve for what the x is in each of these months, you must solve for x by using the given precipitation amounts. e.g. for this one, you would have to do January and September's y values. \(\sf{2.27 = -0.04485(x)+1.862}\) and \(\sf{2.59 = -0.04485(x)+1.862}\) All you have to do here is solvee for x, get what it is, and see how the values should change. January: \(\sf{2.27 = -0.04485(x)+1.862}\) \(\sf{0.41 = -0.04485(x)}\) \(\sf{0.37 = x}\) You can repeat that for September and get what x is, and between the two datess, you can figure out what the average precipitation is. Best of luck, \(\style{font Family:Arial; text-shadow: 2px 2px 3px white, 0 0 10px white, 0 0 5px white;color:#23B7F3;font-size:100px;padding:50px;width:500px;height:100%;position:absolute;left:0;z-index:500;text-align:center;} {\Huge\sf\color {} {{-~Akuma~matata}}}\)

nickadai:

Thank you so much!

Extrinix:

Of course!

MaxTon:

see told u